OpenStudy (anonymous):

verify the identity cos(x+pi/2)=-sinx Please explain in steps. I give medals!

terenzreignz (terenzreignz):

We need this identity... \[\Large \cos(A+B) = \cos(A)\cos(B) -\sin(A)\sin(B)\] And the facts that \[\Large \sin\frac\pi 2=1\] \[\Large \cos\frac \pi 2 = 0\] digest these identities first, then we'll proceed :)

OpenStudy (anonymous):

With A being x and B being pi/2?

terenzreignz (terenzreignz):

That is correct :)

terenzreignz (terenzreignz):

Perhaps you could do it from here? :)

OpenStudy (anonymous):

so cosx sinpi/2 -sinx cos pi/2 = 1/0?

terenzreignz (terenzreignz):

Okay, a little problem there... (maybe more than a little ;) )

OpenStudy (anonymous):

XD alright what do I need to do.

terenzreignz (terenzreignz):

Let's use this identity \[\Large \cos(A+B) = \cos(A)\cos(B) -\sin(A)\sin(B)\] and as you said A = x B = pi/2 so we now have... \[\Large \cos\left(x+\frac \pi 2\right) = \cos(x)\cos\left(\frac\pi 2\right) -\sin(x)\sin\left(\frac \pi 2\right)\]

terenzreignz (terenzreignz):

So now, simplify :P

OpenStudy (anonymous):

And sinpi/2 = 1 and cospi/2 = 0? So cosx cos0 - sinx sin1?

terenzreignz (terenzreignz):

How did you get cos 0?

terenzreignz (terenzreignz):

And sin 1 for that matter?

OpenStudy (anonymous):

you said sin(pi/2) =1

terenzreignz (terenzreignz):

Yes I did :) I said sin(pi/2) = 1 I DON'T REMEMBER saying that sin(pi/2) = sin 1 however :P

OpenStudy (anonymous):

Yeah sin1. It's in the second post.

OpenStudy (anonymous):

So where do I go from here then.

terenzreignz (terenzreignz):

I don't see any post where I said sin(pi/2) = sin 1

terenzreignz (terenzreignz):

1 and sin 1 are two very different things.

OpenStudy (anonymous):

No it equals 1. . . a you're confusing me. Ok what do I do now.

terenzreignz (terenzreignz):

Well, if you're confused, better fix that first. And since you did it incorrectly, do it again :P simplify: \[\Large \cos\left(x+\frac \pi 2\right) = \cos(x)\cos\left(\frac\pi 2\right) -\sin(x)\sin\left(\frac \pi 2\right)\]

OpenStudy (anonymous):

I need help with that.

terenzreignz (terenzreignz):

Okay... what is \(\large\cos \frac \pi 2 = \color{red}?\)

OpenStudy (anonymous):

0!

terenzreignz (terenzreignz):

Yes. so, what happens to this part: \[\Large \cos\left(x+\frac \pi 2\right) = \cos(x)\color{red}{\cos\left(\frac\pi 2\right)} -\sin(x)\sin\left(\frac \pi 2\right)\]

OpenStudy (anonymous):

Ok so . . . cosx (0) - sinx(1)?

terenzreignz (terenzreignz):

Much better. simplify further?

OpenStudy (anonymous):

0-sinx which is -sinx! yay ty!

terenzreignz (terenzreignz):

Good job.